A small nation permits free trade in good X. At the good’s free-trade price of $8, domestic firms supply 6 million units and imports account for 3.2 million units. Recently, the small country has erected trade barriers with the result that imports have fallen to zero, price has risen to $10, and domestic supply has increased to 8 million units. Calculate the change in consumer surplus and producer surplus resulting from the trade barrier. What is the deadweight loss?

Hint: Write your answer to one decimal place.
Answer #1

he free trade price is = 8, Where Quantity supplied =6 Million and Imports =3.2 Million so Quantity demanded = 6+3.2 =9.2 Million

When the trade falls to zero it means that the price is equal to the domestic equilibrium price where the country's imports = 0 therefore,

After trade barriers, equilibrium Price =10

Quantity Supplied=Quantity Demanded =8 Million units

Consumer Surplus in free trade is denoted by the area C+D+E+F

CS decreases by area D+E+F after the trade barrier=

= 10-8* 8 million + 0.5*(9.2 million-8 million)*(10-8)

= 16 million + 1.2 million

= 17.2 million

Producer Surplus in free trade = area P

PS increases by area D after the trade barrier

= [ 05*(8 Million-6 Million)*(10-8)] + [(10-8)*6 Million ]

= 2M+12M = $14 Million

DWL is the area E+F

= 0.5*(8 million-6 million)*(10-8) + 0.5*(9.2 million-8 million)*(10-8)

=2 million+1.2 million

= 3.2 million

Comment